(5x-3x)-2=15. What is the value of X
Answer by waynest(281)

( 5x - 3x ) - 2 = 15
2x - 2 = 15
2x -2 + 2 = 15 + 2
2x = 17
2x/2 = 17/2
x = 17/2 or 8.5
check:
(5(8.5) - 3(8.5)) -2 = 15
(42.5 - 25.5) - 2 = 15
17 -2 = 15
15 = 15